OpenStudy (anonymous):

How much would $150 invested at 8% interest compounded annually be worth after 13 years?

OpenStudy (campbell_st):

use the formula \[ A = P(1 + \frac{r}{100})^n\] A = future value, P = principal, r = interest rate and n = number of compounding periods. in you question you need to find A P = 150, r = 8 and n = 13. Substitute and the evaluate to get the amount.

OpenStudy (anonymous):

can you do the problem.

OpenStudy (campbell_st):

well I'll do this \[A = 150 \times(1 + \frac{8}{100})^{13}\]

OpenStudy (anonymous):

-________________-

OpenStudy (campbell_st):

here is an online calculator to help http://web2.0calc.com/